# OpenS2V-Nexus: A Detailed Benchmark and Million-Scale Dataset for Subject-to-Video Generation

### v3

@@ Abstract
Subject-to-Video (S2V) generation aims to create videos that faithfully incorporate reference content, providing enhanced flexibility in the production of videos. To establish the infrastructure for S2V generation, we propose OpenS2V-Nexus, consisting of (i) OpenS2V-Eval, a fine-grained benchmark, and (ii) OpenS2V-5M, a million-scale dataset. In contrast to existing S2V benchmarks inherited from VBench Vbench that focus on global and coarse-grained assessment of generated videos, OpenS2V-Eval focuses on the model’s ability to generate subject-consistent videos with natural subject appearance and identity fidelity. For these purposes, OpenS2V-Eval introduces 180 prompts from seven major categories of S2V, which incorporate both real and synthetic test data. Furthermore, to accurately align human preferences with S2V benchmarks, we propose three automatic metrics, NexusScore, NaturalScore, and GmeScore, to separately quantify subject consistency, naturalness, and text relevance in generated videos. Building on this, we conduct a comprehensive evaluation of 16 representative S2V models, highlighting their strengths and weaknesses across different content. Moreover, we create the first open-source large-scale S2V generation dataset OpenS2V-5M, which consists of five million high-quality 720P subject-text-video triples. Specifically, we ensure subject-information diversity in our dataset by (1) segmenting subjects and building pairing information via cross-video associations and (2) prompting GPT-Image on raw frames to synthesize multi-view representations. Through OpenS2V-Nexus, we deliver a robust infrastructure to accelerate future S2V generation research.  [^1] The source data and code are publicly available on https://pku-yuangroup.github.io/OpenS2V-Nexus.
@@ 4.1 Data Construction
For Cross-Frame Pairs: since video clips are extracted from long videos, where there exists an inherent temporal and semantic correlation between clips. To capture this, we aggregate clips from the same long video, denoted as $C=\{C_{1},C_{2},\dots,C_{m}\}$, where each $C_{i}$ corresponds to a different segment of the video. The similarity between subjects across these clips is computed using a multimodal retrieval model gme $\mathcal{M}_{\text{retrieval}}$, which computes the similarity score $S(C_{ij},C_{kl})$ for any pair of clips $C_{ij}$ and $C_{kl}$, where $i\neq k$ represents different segments of the video, and $j$ and $l$ represent different subjects:
@@ 5.1 Evaluation Setups
Evaluation Baseline.    We evaluate almost all S2V models, including four closed-source and twelve open ones, including models that support all type of subject (e.g., Vidu vidu , Pika pika , Kling KeLing , VACE vace , Phantom phantom , SkyReels-A2 skyreelsa2 , and HunyuanCustom hunyuancustom ), as well as models that only support human identity (e.g., Hailuo hailuo , ConsisID consisid , Concat-ID concat-id , FantasyID fantasyid , EchoVideo echovideo , VideoMaker videomaker , and ID-Animator ID-Animator ).
@@ 5.2 Comprehensive Analysis
Concat-ID concat-id Open-Source Human-Domain 53.32% 44.13% 31.76% 43.83% 73.67% 66.44%
@@
MAGREF † Ours Human-Domain 52.77% 47.48% 19.19% 48.57% 67.40% 68.11%
MAGREF ‡ Ours Human-Domain 56.12%
(+3.35%)
49.41%
(+1.93%)
23.38%
(+4.19%)
56.51%
(+7.94%)
67.56%
(+0.16%)
69.77%
(1.66%)
@@
Validation of OpenS2V-5M.    Finally, to evaluate the effectiveness and robustness of OpenS2V-5M, we perform qualitative validation based on Wan2.1 wan21 . Specifically, we fine-tune the model using the MAGREF magref method with the Wan2.1 14B weights. Due to limited computational resources, we select 0.5 million data from OpenS2V-5M for training, focusing exclusively on human identity. The results, presented in Figure 9(b) and Table 7, demonstrate that our dataset successfully converts a text-to-video model into a subject-to-video model, thus validating the proposed dataset and its data collection pipeline, especially the Nexus Data plays a crucial role. Since the model is not fully trained, it has not yet achieved optimal performance and is intended for verification purposes only.
@@ Appendix A Related Works: Subject-Consistency Video Generation Models
Diffusion models are widely acknowledged for their remarkable generative capabilities [31, 74, 73, 71, 98, 83, 112, 24], which have significantly advanced the development of subject-consistency generation models [40, 28, 27, 10]. Initially, researchers utilized tuning-based methods to generate consistent image content, such as DreamBooth [76], Lora [32], and Textual Inversion [25]. These methods integrate specific reference content into the training process through fine-tuning existing parameters, adding extra parameters, or modifying text embeddings. Later models, including Magic-Me [64], MotionBooth [102], and DreamVideo [97], extended these approaches to video generation. However, since these methods require training on each new reference content before inference, their practical application is limited. To mitigate the high computational cost, tuning-free methods were introduced. A notable example is IP-Adapter [110], which leverages large datasets to train additional adapters for open-domain subject-consistency generation. However, due to its lower fidelity to human identity, InstantID [91] and PhotoMaker [49] developed human-domain subject-consistency generation models based on this approach. Similar to these image consistency techniques, ID-Animator [30] and ConsisID [113] achieved tuning-free Subject-to-Video (S2V) generation on UNet and DiT, respectively. Nevertheless, these approaches [123, 96, 23, 120] are confined to the human domain, limiting their broader applicability. Recent works, such as Phantom [57], VACE [42], and SkyReels-A2 [22], have demonstrated the ability to generate consistent multi-subject videos in the open domain [51, 13, 37], gradually narrowing the gap with commercial S2V models [45, 46, 86, 5]. However, a unified and comprehensive benchmark to assess the strengths and weaknesses of these models remains absent, and the lack of publicly released training data impedes further progress in this field. Therefore, we introduce OpenS2V-Eval and OpenS2V-5M, aimed at bridging this gap.
@@ C.3 Further Verification on OpenS2V-5M
Due to limited space in the main text, we provide additional qualitative analysis of MAGREF ‡here, with results shown in Figure 16. It can be observed that MAGREF ‡is capable of generating high-quality videos, thereby validating the effectiveness of the proposed OpenS2V-5M.
@@ D.2 Details of Evaluation Models
Phantom    Model Details. Phantom [57] is a video generation model based on DiT that extracts reference image information using both CLIP and VAE, and employs a windowed attention mechanism to reduce computational overhead, enabling open-domain subject-to-video generation. It includes three model weights: Phantom-Seaweed, Phantom-Wan-1.3B, and Phantom-Wan-14, but only Phantom-Wan-1.3B is publicly released. The training data come from panda70M [12], subject200k [14], OmniGen [104], and internal datasets, totaling over 10 million samples. Implementation Setups. We use the officially released Phantom-Wan-1.3B code and model, maintaining the original settings. We generate 5-second (81-frame) videos at a resolution of $832\times 480$ and a 16 fps.
SkyReels-A2    Model Details. SkyReels-A2 [22] is a model fine-tuned based on Wan2.1 [88], employing an approach similar to Phantom. It utilizes a dual-stream architecture to enhance the model’s response to reference images and textual prompts, enabling open-domain subject-to-video generation. There are four variants in total: A2-Wan2.1-14B-Preview, A2-Wan2.1-14B, A2-Wan2.1-14B-Pro, and A2-Wan2.1-14B-Infinity, but only A2-Wan2.1-14B-Preview has been open-sourced. The training data comes from 2 million high-quality subject-text-video triples collected internally. Implementation Setups. We use the officially released A2-Wan2.1-14B-Preview code and model, maintaining the original settings. Videos are generated with a spatial resolution of $832\times 480$ and a frame rate of 16 fps, resulting in a duration of 5 seconds (81 frames).
@@
Concat-ID    Model Details. Concat-ID [123] is a model fine-tuned based on CogVideoX [113]. It concatenates image features with video latents along the token dimension, thereby avoiding the issue of blurry initial frames. It only supports input of a single face image. The training data is processed from internally collected data, with a dataset size of approximately 1.3 million. Implementation Setups. We use the officially released Concat-ID code and model, maintaining the original settings. Videos are generated with a spatial resolution of $720\times 480$ and a frame rate of 8 fps, resulting in a duration of 6 seconds (49 frames).
@@ E.3 Potential Harms Caused by the Research Process
The subject images of OpenS2V-Eval are derived from three open-source datasets—ConsisID [113], A2-Bench [22], and DreamBench [70]—that adhere to the Apache license, as well as from three video platforms—Pexels, MixKit, and PixaBay—that operate under the Creative Commons Zero (CC0) license. The video data in OpenS2V-5M originates from the Open-Sora Plan [52], with some content licensed under Creative Commons Attribution 4.0 (CC BY 4.0) and others under the Royalty-Free (RF) license. The licensing information for these data is explicitly stated on their respective platforms. The CC0 license designates content as public domain, permitting unrestricted use without additional permissions or authorizations. For CC BY 4.0-licensed videos from the Open-Sora Plan [52], video IDs are included in the metadata to mitigate potential contractual disputes. For RF-licensed videos, we are working to resolve intellectual property issues. In total, approximately 4 million data will be made available as open source. The collected data is organized into seven categories, with contributions from global sources. This diversity ensures that OpenS2V-Eval and OpenS2V-5M are fully representative. The MAGREF model [19] fine-tuned on our dataset demonstrated no significant content bias. Furthermore, video content has been filtered to exclude NSFW material based on subtitle detection. Due to the presence of videos containing identifiable individuals, access to OpenS2V-Nexus is restricted to academic use only, with contact information provided on the https://pku-yuangroup.github.io/OpenS2V-Nexus to ensure the security of personal identity data.

### v4

@@ Abstract
Subject-to-Video (S2V) generation aims to create videos that faithfully incorporate reference content, providing enhanced flexibility in the production of videos. To establish the infrastructure for S2V generation, we propose OpenS2V-Nexus, consisting of (i) OpenS2V-Eval, a fine-grained benchmark, and (ii) OpenS2V-5M, a million-scale dataset. In contrast to existing S2V benchmarks inherited from VBench Vbench that focus on global and coarse-grained assessment of generated videos, OpenS2V-Eval focuses on the model’s ability to generate subject-consistent videos with natural subject appearance and identity fidelity. For these purposes, OpenS2V-Eval introduces 180 prompts from seven major categories of S2V, which incorporate both real and synthetic test data. Furthermore, to accurately align human preferences with S2V benchmarks, we propose three automatic metrics, NexusScore, NaturalScore, and GmeScore, to separately quantify subject consistency, naturalness, and text relevance in generated videos. Building on this, we conduct a comprehensive evaluation of 18 representative S2V models, highlighting their strengths and weaknesses across different content. Moreover, we create the first open-source large-scale S2V generation dataset OpenS2V-5M, which consists of five million high-quality 720P subject-text-video triples. Specifically, we ensure subject-information diversity in our dataset by (1) segmenting subjects and building pairing information via cross-video associations and (2) prompting GPT-Image on raw frames to synthesize multi-view representations. Through OpenS2V-Nexus, we deliver a robust infrastructure to accelerate future S2V generation research.  [^1] The source data and code are publicly available on https://pku-yuangroup.github.io/OpenS2V-Nexus.
@@ 4.1 Data Construction
For Cross-Frame Pairs: since clips are split from long videos, where there exists an inherent temporal and semantic correlation between clips concat-id . To capture this, we aggregate clips from the same long video, denoted as $C=\{C_{1},C_{2},\dots,C_{m}\}$, where each $C_{i}$ corresponds to a different segment of the video. The similarity between subjects across these clips is computed using a multimodal retrieval model gme $\mathcal{M}_{\text{retrieval}}$, which computes the similarity score $S(C_{ij},C_{kl})$ for any pair of clips $C_{ij}$ and $C_{kl}$, where $i\neq k$ represents different segments of the video, and $j$ and $l$ represent different subjects:
@@ 5.1 Evaluation Setups
Evaluation Baseline.    We evaluate almost all S2V models, including four closed-source and fourteen open ones, including models that support all type of subject (e.g., Vidu vidu , Pika pika , Kling KeLing , VACE vace , Phantom phantom , SkyReels-A2 skyreelsa2 , and HunyuanCustom hunyuancustom ), as well as models that only support human identity (e.g., Hailuo hailuo , ConsisID consisid , Concat-ID concat-id , FantasyID fantasyid , EchoVideo echovideo , VideoMaker videomaker , and ID-Animator ID-Animator ).
@@ 5.2 Comprehensive Analysis
Phantom-14B phantom Open-Source Open-Domain 58.69% 49.14% 41.24% 55.02% 72.55% 68.33%
@@
Concat-ID-CogVideoX concat-id Open-Source Human-Domain 53.32% 44.13% 31.76% 43.83% 73.67% 66.44%
Concat-ID-Wan-AdaLN concat-id Open-Source Human-Domain 53.18% 43.13% 17.19% 50.05% 71.90% 69.44%
@@
Ours † - Human-Domain 51.67% 41.30% 20.83% 47.64% 72.12% 65.42%
Ours ‡ - Human-Domain 52.97%
(+1.30%)
41.86%
(+0.56%)
22.77%
(+1.94%)
49.51%
(+1.87%)
72.35%
(+0.23%)
66.80%
(+1.38%)
@@
Validation of OpenS2V-5M.    Finally, to evaluate the effectiveness and robustness of OpenS2V-5M, we fine-tune a model initialized with Wan2.1 1.3B weights wan21 using the ConsisID method consisid , employing only MSE loss and omitting mask loss. Given computational constraints, we randomly use 300k samples from OpenS2V-5M, focusing solely on single human identity during training. The results, presented in Figure 9(b) and Table 7, demonstrate that our dataset successfully converts a text-to-video model into a subject-to-video model, thus validating the proposed dataset and its data collection pipeline, especially the Nexus Data plays a crucial role. Since the model is not fully trained, it has not yet achieved optimal performance and is intended for verification purposes only.
@@ Appendix A Related Works: Subject-Consistency Video Generation Models
Diffusion models are widely acknowledged for their remarkable generative capabilities [75, 74, 72, 63, 64, 62, 84, 112, 23], which have significantly advanced the development of subject-consistency generation models [38, 27, 26, 10]. Initially, researchers utilized tuning-based methods to generate consistent image content, such as DreamBooth [77], Lora [30], and Textual Inversion [24]. These methods integrate specific reference content into the training process through fine-tuning existing parameters, adding extra parameters, or modifying text embeddings. Later models, including Magic-Me [65], MotionBooth [102], and DreamVideo [98], extended these approaches to video generation. However, since these methods require training on each new reference content before inference, their practical application is limited. To mitigate the high computational cost, tuning-free methods were introduced. A notable example is IP-Adapter [110], which leverages large datasets to train additional adapters for open-domain subject-consistency generation. However, due to its lower fidelity to human identity, InstantID [92] and PhotoMaker [47] developed human-domain subject-consistency generation models based on this approach. Similar to these image consistency techniques, ID-Animator [29] and ConsisID [113] achieved tuning-free Subject-to-Video (S2V) generation on UNet and DiT, respectively. Nevertheless, these approaches [123, 97, 22, 120] are confined to the human domain, limiting their broader applicability. Recent works, such as Phantom [55], VACE [40], and SkyReels-A2 [21], have demonstrated the ability to generate consistent multi-subject videos in the open domain [49, 13, 35], gradually narrowing the gap with commercial S2V models [43, 44, 87, 5]. However, a unified and comprehensive benchmark to assess the strengths and weaknesses of these models remains absent, and the lack of publicly released training data impedes further progress in this field. Therefore, we introduce OpenS2V-Eval and OpenS2V-5M, aimed at bridging this gap.
@@ C.3 Further Verification on OpenS2V-5M
[…]
@@ D.2 Details of Evaluation Models
Phantom    Model Details. Phantom [55] is a video generation model based on DiT that extracts reference image information using both CLIP and VAE, and employs a windowed attention mechanism to reduce computational overhead, enabling open-domain subject-to-video generation. It includes three model weights: Phantom-Seaweed, Phantom-Wan-1.3B, and Phantom-Wan-14, but only Phantom-Wan-1.3B&14B are publicly released. The training data come from panda70M [12], subject200k [14], OmniGen [104], and internal datasets, totaling over 10 million samples. Implementation Setups. We use the officially released Phantom-Wan code and model, maintaining the original settings. We generate 5-second (81-frame) videos at a resolution of $832\times 480$ and a 16 fps.
SkyReels-A2    Model Details. SkyReels-A2 [21] is a model fine-tuned based on Wan2.1 [89], employing an approach similar to Phantom. It utilizes a dual-stream architecture to enhance the model’s response to reference images and textual prompts, enabling open-domain subject-to-video generation. There are four variants in total: A2-Wan2.1-14B-Preview, A2-Wan2.1-14B, A2-Wan2.1-14B-Pro, and A2-Wan2.1-14B-Infinity, but only A2-Wan2.1-14B-Preview has been open-sourced. The training data comes from 2 million high-quality subject-text-video triples collected internally. Implementation Setups. We use the officially released SkyReels-A2-Wan2.1-14B-Preview code and model, maintaining the original settings. Videos are generated with a spatial resolution of $832\times 480$ and a frame rate of 16 fps, resulting in a duration of 5 seconds (81 frames).
@@
Concat-ID    Model Details. Concat-ID [123] is a model fine-tuned based on CogVideoX [113] and Wan2.1 [89]. It concatenates image features with video latents along the token dimension, thereby avoiding the issue of blurry initial frames. It only supports input of a single face image. The training data is processed from internally collected data, with a dataset size of approximately 1.3 million. Implementation Setups. We use the officially released Concat-ID code and model, maintaining the original settings. For CogVideoX version, videos are generated with a spatial resolution of $720\times 480$ and a frame rate of 8 fps, resulting in a duration of 6 seconds (49 frames). For Wan-AdaLN version, videos are generated with a spatial resolution of $832\times 480$ and a frame rate of 16 fps, resulting in a duration of 5 seconds (81 frames).
@@ E.3 Potential Harms Caused by the Research Process
The subject images of OpenS2V-Eval are derived from three open-source datasets—ConsisID [113], A2-Bench [21], and DreamBench [71]—that adhere to the Apache license, as well as from three video platforms—Pexels, MixKit, and PixaBay—that operate under the Creative Commons Zero (CC0) license. The video data in OpenS2V-5M originates from the Open-Sora Plan [50], with some content licensed under Creative Commons Attribution 4.0 (CC BY 4.0) and others under the Royalty-Free (RF) license. The licensing information for these data is explicitly stated on their respective platforms. The CC0 license designates content as public domain, permitting unrestricted use without additional permissions or authorizations. For CC BY 4.0-licensed videos from the Open-Sora Plan [50], video IDs are included in the metadata to mitigate potential contractual disputes. For RF-licensed videos, we are working to resolve intellectual property issues. In total, approximately 4 million data will be made available as open source. The collected data is organized into seven categories, with contributions from global sources. This diversity ensures that OpenS2V-Eval and OpenS2V-5M are fully representative. The ConsisID model [113] fine-tuned on our dataset demonstrated no significant content bias. Furthermore, video content has been filtered to exclude NSFW material based on subtitle detection. Due to the presence of videos containing identifiable individuals, access to OpenS2V-Nexus is restricted to academic use only, with contact information provided on the https://pku-yuangroup.github.io/OpenS2V-Nexus to ensure the security of personal identity data.